Given that R1=3i-2j+k, R2=2i-4j-3k, R3=-i+2j+2k.
Find the magnitude of

A=2r1-3r2-5r3
1
Expert's answer
2021-07-09T08:31:48-0400
"\vec{A}=2(3,-2,1)-3(2,-4,-3)-5(-1,2,2)\\\vec{A}=(5,-2,1)"